# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Model is not a General Substitute for GPT-4

### v3

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-sourced models, especially GPT-4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this work, we conduct an empirical study of judge models. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness, aspect-specific evaluation, and scalability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations. Finally, we introduce a integrated method, leveraging GPT-4 to compensate for the limitations and improve the fine-tuned judges. Experiment results show our method achieves accuracy on par with GPT-4 with only 50% of the API expense. [^1] Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn significant attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them in a scalable and comprehensive manner Qin et al. (2023). Some research has proposed LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, proprietary LLMs can provide an accurate evaluation with a high agreement rate with human evaluators.
Figure 1: The general training and inference procedure of fine-tuned judge models. The training data generally comprises triplets with instruction, response and evaluation result, which are fed to the open-sourced foundation model (such as LLaMA) to create a judge model.
@@
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2023b); Wang et al. (2024); Ke et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
With the prevalence of fine-tuned judge models, in this paper, we conduct an empirical study of their evaluation capabilities. Our study encompasses a comprehensive comparison across various benchmarks and dimensions. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced models in the following aspects:
  • The fine-tuned judge model is biased towards superficial quality;
  • The fine-tuned judge model can not benefit from prompt engineering;
We argue that these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data.
In light of the empirical findings, we propose an integrated evaluation method, CascadedEval, to compensate for the limitations with the help of a proprietary model. CascadedEval is based on the confidence indicator derived from softmax probability distribution to assess the confidence of the fine-tuned judge. After that, samples with lower confidence are re-evaluated by a proprietary model, such as GPT-4. Experiment results show our method achieves accuracy on par with GPT-4 on both in-domain and out-of-domain testsets, with only 50% of the API expense.
Our contributions can be summarized as follows:
  • We reveal several limitations of fine-tuned judge models for LLM evaluation by conducting a comprehensive empirical study;
  • We propose an integrated evaluation method by leveraging GPT-4 to compensate for the limitations of fine-tuned judges;
  • To the best of our knowledge, this is the first systematic study of the limitations of fine-tuned judge models for LLM evaluation;
2 Related Work
2.1 LLM-as-a-Judge
Benchmark evaluation has long been the default standard in the study of traditional pre-trained models Wang et al. (2018, 2019). Therefore, after the emergence of LLMs, numerous efforts have been made to construct benchmark testsets that cover a variety of tasks for evaluation Srivastava et al. (2023); Zhong et al. (2024). However, questions in these benchmarks are mostly formalized as multi-choice selection or completion, which can not fully showcase the models’ generation capability.
Consequently, recent research has introduced the concept of LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, Li et al. (2023c) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT-4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT-4. The GPT-4-based evaluator is proven to be high accurate compared with professional human evaluators and presents even better consistency and stability compared with human. However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility.
@@ 2.1 LLM-as-a-Judge
(Zhu et al., 2023b) Vicuna Instruct Datasets
@@
(Li et al., 2023a) LLaMA2-chat Preference Datasets
@@
(Kim et al., 2023) LLaMA2-chat GPT-4 Generated GPT-4 Generated GPT-4 Pointwise Grading GPT-4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study. All four models are open-source, with their training and test data also publicly released.
2.2 Finetuned LLM as a Judge
In response to the concerns posed by the close-sourced models, follow-up works suggest fine-tuning language models locally for evaluations. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna Chiang et al. (2023) as a scalable judge model. Auto-J Li et al. (2023a) constructs judgment data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria, constructs a feedback dataset based on GPT-4, and finetunes a fine-grained judge model.
Despite relying on models with 7B or 13B parameters, these fine-tuned judge models all achieve comparable accuracy with GPT-3.5 or GPT-4 on their respective testsets. However, the evaluation is mostly conducted on the in-domain testsets constructed similarly with the training data. A thorough examination of the evaluation capability of the fine-tuned judges is in urgent demand.
@@ 3 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models as listed in Table 1.
3.1 Preliminary: Fine-tuned LLM Judge
The typical process for finetuning a judge model consists of the following three steps:
Step 1: Data Collection. The training data generally comprises three components: instructions, responses and evaluations. The instructions are typically obtained from instruction datasets, with the responses generated by various representative models, and the evaluations can be derived from either GPT-4 or human annotation.
Step 2: Prompt Designing. The prompt template can be structured in various ways depending on the evaluation scheme, such as pairwise selection (which aims to select the better one from a pair of responses to the instruction), and pointwise grading (which aims to assign a score to a single response based on the instruction), etc.
Step 3: Model Fine-tuning. Using the designed prompt and collected data, the training process of the judge model typically follows the instruction fine-tuning paradigm Ouyang et al. (2022). The model is fed with a instruction alongside response(s) to generate output, which includes evaluation results and possibly explanations.
After that, the fine-tuned judge can be adopted in evaluating the output of LLM. While these models are able to achieve superior performance on their self-designed testsets, we reveal that there exists several limitations about the evaluation capabilities of the fine-tuned judges.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
JudgeLM-7B 82.39 72.97 68.17 65.18 45.3 0.398 0.384
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 0.417 0.386
Auto-J-13B 77.79 62.64 72.17 64.10 53.6 0.614 0.591
Prometheus-13B 24.58 23.39 29.03 27.92 16.2 0.864 0.869
+grade-twice 54.24 50.04 45.25 43.58 47.8
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106 84.24 72.83 75.78 71.51 56.9 0.742 0.743
Table 2: Results of evaluators on pairwise selection. Notice Prometheus can be transformed for pairwise selection by grading each answers and compare the scores, therefore we release both results with and without grade-twice.
3.2 Limitations of Fine-tuned Judges
3.2.1 Constrained by Evaluation Scheme
Model MT-Bench
accuracy precision recall F1
JudgeLM-7B 48.7 52.0 49.7 48.7
PandaLM-7B 55.2 52.6 49.4 46.8
Auto-J-13B 51.7 50.2 46.8 43.7
Prometheus-13B 53.2 49.6 48.4 47.1
GPT-4-1106 66.9 63.8 62.2 61.9
Table 3: Results of evaluators on multi-turn evaluation.
@@ 3.2.1 Constrained by Evaluation Scheme
One of the most appealing attributes of LLMs is their generalization ability, enabling them to execute various tasks defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their evaluation capability under the scheme defined by others. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets [^2] We make minimal modifications to the predefined prompts to adapt the model to different schemes. For detailed prompts please refer to Appendix A.2..
As shown in Table 2, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained, the evaluation performance would drop by a large margin. For example, using a pairwise model for pointwise grading, or using a pointwise model for pairwise selection, would all lead to catastrophic performance degradation. On the contrary, close-sourced models such as GPT-3.5 or GPT-4 consistently exhibit superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset\@footnotemark. As shown in Table 3, while the four models are all trained for single-turn evaluation, they underperform GPT-4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lost their generalizability.
3.2.2 Biased Towards Superficial Quality
Recently, there has been a lot of research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). Subsequently, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.). We evaluate the judge models on LLMBar [^3] The detailed prompts are present in Appendix A.2. as shown in Table 4.
As can be seen, the fine-tuned judge models perform poorly on adversarial testsets, even worse than random-guess. This notifies that they are severely biased toward superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference for incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This notifies that the superior performance of fine-tuned judges on the in-domain testsets may rely on spurious features Niven and Kao (2019), instead of really differentiating response quality.
Model LLMBar
Natu. Neig. GPTI. GPTO. Manu.
JudgeLM-7B 62.0 23.1 26.1 46.8 28.3
PandaLM-7B 59.0 16.5 21.7 42.6 26.1
Auto-J-13B 70.0 20.9 21.7 46.8 23.9
Prometheus-7B 53.0 22.4 17.4 27.7 32.6
DeBERTa 62.0 26.9 42.4 55.3 34.8
GPT-4-1106 93.5 64.2 76.6 76.6 75.0
Table 4: Accuracy of evaluators on bias evaluation. Neig., GPTI., GPTO., and Manu. are the four adversarial test sets designed to quantify the bias.
@@ 3.2.2 Biased Towards Superficial Quality
Table 5: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. As HaluEval and ToxicChat are both binary classifications, we apply Auto-J and Prometheus with pointwise grading and conduct a grid search to determine the classification threshold. On the other hand, as SALAD-Bench is a pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM, and Auto-J to select a better response.
We also fine-tune a DeBERTa-based judge model in a classification style on LLMBar (please refer to Section 3.3 for details). It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluators may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation.
3.2.3 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as helpfulness, safety, etc., and sometimes we are particularly interested in a specific aspect. While previous work primarily assess the judge models from a general perspective, in this part, we would like to assess their evaluation capability on fine-grained aspects. We select the following three datasets:
  1. 1.
    HaluEval Li et al. (2023b): This dataset is used for factuality evaluation. It consists of three domains: Question-Answering, Summary, and Dialogue.
  2. 2.
    ToxicChat Lin et al. (2023): This dataset is used for toxicity evaluation.
  3. 3.
    SALAD-Bench Li et al. (2024): This dataset is used for safety evaluation.
We validate both close-sourced and fine-tuned judges on the three datasets [^4] We make minimal modification to the prompts to adapt them to the specific aspects, as detailed in Appendix A.2.. As can be seen from Table 5, the fine-tuned judges fall far behind the close-sourced judges on all fine-grained aspects. It deserves to notice that while Prometheus is designed for fine-grained evaluation, it obtains an inferior performance on both benchmarks, which notifies that it failed to learn the correlation between fine-grained aspects and evaluation results.
3.2.4 No Benefits from Prompt Engineering
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies and validate their effectiveness for fine-tuned judge models:
@@ 3.2.4 No Benefits from Prompt Engineering
Table 6: Results of evaluators with ICL and CoT. Increased results are in bold while decreased results are in grey. We did not apply GPT-4 on JudgeLM-test as the annotation of JudgeLM-test is conducted with GPT-4 without ICL and CoT. We only apply ICL on Auto-J as the original prompt of Auto-J comprises CoT.
  1. 1.
    In-context Learning (ICL) Dong et al. (2023): where task demonstrations are integrated into the prompt as illustration. In our work, we randomly select 2-4 demonstrations from the training set based on the max-context length.
  2. 2.
    Chain-of-thought (CoT) Wei et al. (2022): where the input prompt is structured in a way that mimics human reasoning. In our work, the judge model is forced to generate the explanation first, then provide a final judgment.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
Released Models 82.39 72.97 68.97 60.95 53.6 0.864 0.869
Vicuna-generation 82.44 71.77 72.37 60.78 47.6 0.826 0.815
Vicuna-classification 82.16 70.07 70.87 60.34 46.8 0.846 0.831
DeBERTa-classification 81.30 68.34 72.27 51.75 31.7 0.835 0.813
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106-preview 84.24 72.83 75.78 71.51 56.9 0.742 0.743

### v4

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have fine-tuned judge models based on open-source LLMs for evaluation. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this work, we conduct an empirical study of LLM-as-a-Judge. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness and adaptability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations1.
footnotetext: * Equal contribution. ✉ Corresponding Author.footnotetext: 1Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge.
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn significant attention Liang et al. (2022); Chang et al. (2023); He et al. (2024); Gu et al. (2025); He et al. (2025). Some research has proposed LLM-as-a-Judge Li et al. (2023b); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, proprietary LLMs can provide an accurate evaluation with high agreement with human evaluators.
Figure 1: The general training and inference procedure of fine-tuned judge models.
@@
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2024); Wang et al. (2024); Ke et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
@@
(Zhu et al., 2024) Vicuna Instruct Datasets
@@
(Li et al., 2024a) LLaMA2-chat Preference Datasets
@@
(Kim et al., 2024) LLaMA2-chat GPT-4 Generated GPT-4 Generated GPT-4 Pointwise Grading GPT-4
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test MT-Bench
accuracy F1 accuracy F1 agreement PCC-ind PCC-ood accuracy F1
JudgeLM-7B 82.39 72.97 68.17 65.18 45.3 0.398 0.384 48.7 48.7
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 0.417 0.386 55.2 46.8
Auto-J-13B 77.79 62.64 72.17 64.10 53.6 0.614 0.591 51.7 43.7
Prometheus-13B 24.58 23.39 29.03 27.92 16.2 0.864 0.869 53.2 47.1
+grade-twice 54.24 50.04 45.25 43.58 47.8
Deepseek-V3 79.23 68.27 75.97 71.25 57.0 0.734 0.741
GPT-4-mini 79.17 68.31 76.57 71.79 57.4 0.707 0.705
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106 84.24 72.83 75.78 71.51 56.9 0.742 0.743 66.9 61.9
Table 2: Results of evaluators on different evaluation schemes. Notice JudgeLM-test, PandaLM-test, Auto-J-test are pairwise selection, Prometheus-test is pointwise grading, and MT-Bench is multi-turn evaluation.
In this paper, we conduct an empirical study for the evaluation capability of judge models. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced proprietary models:
  • The fine-tuned judge model is biased towards superficial quality;
  • The fine-tuned judge model can not benefit from prompting strategies;
We argue that these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data. To draw a conclusion, the fine-tuned judge model cannot serve as a general substitute for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging them for evaluation in real applications, watching for the overlap between the evaluation scenario and the fine-tuning process.
@@ 2 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models in Table 1111We make minimal change to the predefined prompts to adapt the judge model to different schemes. Please refer to Appendix A.2 for detailed implementations., and reveal there exist several limitations about their evaluation capabilities.
2.1 Constrained by Evaluation Scheme
@@
One of the most appealing attributes of LLMs is their generalization ability, enabling them to execute various tasks defined by various instructions Zhu et al. (2023). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, chain-of-thought evaluation, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their capability on uncovered schemes. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets. We also validate the models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset.
@@
Table 3: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. For more details please refer to A.2.
Model LLMBar
Natu. Neig. GPTI. GPTO. Manu.
JudgeLM-7B 62.0 23.1 26.1 46.8 28.3
PandaLM-7B 59.0 16.5 21.7 42.6 26.1
Auto-J-13B 70.0 20.9 21.7 46.8 23.9
Prometheus-7B 53.0 22.4 17.4 27.7 32.6
GPT-4-1106 93.5 64.2 76.6 76.6 75.0
Table 4: Accuracy of evaluators on bias evaluation.
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement PCC-ind PCC-ood
Released Models 82.39 72.97 68.97 60.95 53.6 0.864 0.869
Vicuna-generation 82.44 71.77 72.37 60.78 47.6 0.826 0.815
Vicuna-classification 82.16 70.07 70.87 60.34 46.8 0.846 0.831
DeBERTa-classification 81.30 68.34 72.27 51.75 31.7 0.835 0.813
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106-preview 84.24 72.83 75.78 71.51 56.9 0.742 0.743
As shown in Table 2, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained, the evaluation performance would drop by a large margin. On the contrary, close-sourced proprietary models such as GPT-3.5 or GPT-4 consistently exhibit superior performance across various evaluation schemes.
Model Method JudgeLM-test PandaLM-test Auto-J-test Salad-bench
accuracy F1 acuracy F1 agreement accuracy F1
JudgeLM-7B w/o CoT 82.74 72.64 72.67 69.32 44.25 85.57 57.35
w/ original CoT 82.26 67.41 70.77 64.24 41.45 75.15 50.59
w/ o1 CoT 77.96 65.43 66.46 62.95 37.90 72.39 47.91
Table 6: Comparison of different CoT sources on JudgeLM-7B.
@@
Table 7: Results of evaluators with ICL and CoT. We did not apply GPT-4 on JudgeLM-test as the annotation of JudgeLM-test is conducted with GPT-4 without ICL and CoT. We only apply ICL on Auto-J as the original prompt of Auto-J comprises CoT.
2.2 Biased Towards Superficial Quality
Recently, there has been a lot of research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). Subsequently, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises four adversarial testsets (Neig., Manu., GPTO., GPTI.) with paired outputs of a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.).
We evaluate the judge models on LLMBar. As shown in Table 4, the fine-tuned judge models perform poorly on adversarial testsets, even worse than random-guess. This notifies that they are severely biased toward superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference for incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on LLMBar.
2.3 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as helpfulness, safety, etc. In this part, we would like to assess the evaluation capability of judge models on fine-grained aspects, based on the following datasets: 1) HaluEval Li et al. (2023a): for factuality evaluation; 2) ToxicChat Lin et al. (2023): for toxicity evaluation; 3) SALAD-Bench Li et al. (2024b): for safety evaluation.
As can be seen from Table 3, the fine-tuned judges fall far behind on all fine-grained aspects. It deserves to notice that while Prometheus is designed for fine-grained evaluation, it obtains an inferior performance on both benchmarks, which notifies that it failed to learn the correlation between fine-grained aspects and evaluation results.
2.4 Can not Benefit from CoT and ICL
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies, namely In-context Learning (ICL) Dong et al. (2023) and Chain-of-Thought Prompting (CoT) Wei et al. (2022), to further improve the evaluation capability of the judge models.